Tohban Report 2017-03-29
From RHESSI Wiki
|Start Date:||22 March 2017|
|End Date:||29 March 2017|
|List all reports|
There are 4 named AR on disk. The most active of them had a few C-class flares in the past few days, its still East of disk center, so we can hopefully catch some more activity from it in the coming week. We caught the two C-5's on March 28
How many GOES flares occurred?
Flares above B, C, M, X class were 23 4 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 11 2 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 8 0 0 0
There were RHESSI flares/GOES flares 33 / 27 over the time range 22-Mar-17 28-Mar-17
Decimation was normal/vigorous for the full time period, with the SSR at 0% at the end of each day's contacts.
Detectors 3 and 6 are recording events. The cold tips are 127.1 and 125.5 K as of 29-Mar-2017. The cold plates, 1 and 2 are 153.3 and 151.9 K. Cryocooler power is steady at 75W. See:
for a summary plot of recent temperatures, efficiencies and accelerometer values.
No data gaps over the past week.
Detectors 3 and 6 are powered on, and are recording events. No threshold or HV changes.
There are counts coming in the maximum energy channel on D6. Its a "new" phenomenon, and we're not sure where its coming from or why. Don't think its damage, no high leakage current or resets. Next Tohban keep and eye on it and maybe try some tests to see where it comes from. Adjusting HV was suggested.
Battery has been doing well this week, end of night voltages of 26.2 - 26.3 V, no increase in pressure or temperature profiles.
|HLAT Decimation||Rear decimation|
|Night time data (fronts)||plus/minus 4 minutes from daylight|
|Night time data (rears)||plus/minus 4 minutes from daylight|
|Require extra passes?||No|
|Requirement for moving pointer?||No|